OpenStudy (anonymous):

Find the payment necessary to amortize the loan. Do not round until the final answer. Then, round to the nearest cent. $13,100; 12% compounded monthly; 48 monthly payments.

OpenStudy (anonymous):

PMT = PVr/(1-(1+r)^-n) = 13,100*0.01/(1-1.01^-48) = = $344.97

OpenStudy (anonymous):

you round your equation answers than you have your answer $344.97

OpenStudy (anonymous):

Is this the same? John bought a new computer for $1,650. He paid a $160 down payment and financed the rest for 1 year at an interest rate of 7%. Find the total interest paid on the given amortized loan assuming that John makes monthly payments. Do not round until the final answer. Then, round to the nearest cent.

OpenStudy (anonymous):

I replied to this question! :) You can see it in the questions you asked! But it's a bit the same.